Chris has a miniature model of the boat that he is building. The length of the model is 4.5 inches and the height is 2 inches. If the length of the actual boat is 18 feet, what is the height?

Answer by Alan3354(69443)

Chris has a miniature model of the boat that he is building. The length of the model is 4.5 inches and the height is 2 inches. If the length of the actual boat is 18 feet, what is the height?
------------------
The scale is 4.5:18*12 = 4.5:216 = 1:48
---
Height = 2*48 inches = 8 feet